OpenStudy (mythorius):

The lengths of the sides of a triangle are 6, 7, and 9. Find the longest side of a similar triangle whose shortest side is 4. 2√6 6 13 1/2 @haleyelizabeth2017

OpenStudy (mythorius):

Six?

OpenStudy (haleyelizabeth2017):

\[\frac{4}{6}=\frac{x}{9}\]\[9*4/6=?\]

OpenStudy (haleyelizabeth2017):

yeh

OpenStudy (mythorius):

Lmao that bot is here haha

OpenStudy (haleyelizabeth2017):

bot?

OpenStudy (mythorius):

Look at the people in here

OpenStudy (haleyelizabeth2017):

oh, the ghost

OpenStudy (mythorius):

Yah

OpenStudy (haleyelizabeth2017):

LOL that's just TSO stalking again ;) x'D